# Symplectic Neural Flows for Modeling and Discovery

### v1

@@ Acknowledgements
Appendix A The classical mechanics of non-conservative systems: Damped Harmonic Oscillator
@@
where $L$ describes the conservative contribution to the full non-conservative Lagrangian $\Lambda$, and $K$ is an antisymmetric function under exchange $a\to b$, that couples the variables together and accounts for all the non-conservative contribution to the system dynamics. Notice that, in conservative Hamiltonian systems, the function $K$ would vanish. Although, $K$ also vanishes in the physical limit (PL) [23], that is, in the linear subspace $\{(q_{a},q_{b},p_{a},p_{b})\in\mathbb{R}^{4d}:\,\,q_{a}=q_{b},\text{ and }p_{a}=p_{b}\}$, its derivatives do not necessarily vanish in the physical limit. This is important, since the derivatives of $K$ model the dissipative behavior of the system, as we will see below.
The stationarity of the action in Equation 27 under the variations $q_{I}(t,\epsilon)=q_{I}(t,0)+\epsilon\eta_{I}(t)$ , that is $[\mathrm{d}S[q_{I}]/\mathrm{d}\epsilon]_{\epsilon=0}=0$ for all $\eta_{I}(t_{i})=\eta_{I}(t_{f})=0$, $I=\{a,b\}$, corresponds to the conditions
@@
[…]
@@
We remark that $\pi_{a}=-\pi_{b}$ can be inferred e.g. in the case of a damped harmonic oscillator below by noticing that $\pi_{a}+\pi_{b}=p_{a}-p_{b}-\lambda(q_{a}-q_{b})=0$, where $p_{a}$ and $p_{b}$ are the conservative momenta derived from the conservative Lagrangian as $p_{I}=\partial_{\dot{q}_{I}}L(q_{I},\dot{q}_{I})$, $I\in\{a,b\}$.
@@
$\begin{bmatrix}\dot{q}_{a}(t)\\ \dot{q}_{b}(t)\\ \dot{p}_{1}(t)\\ \dot{p}_{2}(t)\end{bmatrix}=\begin{bmatrix}\frac{\pi_{a}(t)}{m}+\frac{\lambda}{2m}(q_{a}(t)-q_{b}(t))\\ -\frac{\pi_{b}(t)}{m}-\frac{\lambda}{2m}(q_{a}(t)-q_{b}(t))\\ -\frac{\lambda}{2m}(\pi_{a}(t)-\pi_{b}(t))-kq_{a}(t)\\ \frac{\lambda}{2m}(\pi_{a}(t)-\pi_{b}(t))+kq_{b}(t)\end{bmatrix}.$ (35)
Notice that the physical limit in Equation 31 is an invariant submanifold of $\mathbb{R}^{4d}$ with respect to the dynamics, since $\dot{\pi}_{a}+\dot{\pi}_{b}=0$ whenever $q_{a}=q_{b}$, and also $\dot{q}_{a}-\dot{q}_{b}=0$. To ensure that our neural networks preserve this subspace as well, we apply a projection step at the end of the network, which projects over the subspace in Equation 31 and is defined as
@@
We remark that this projection turns the SympFlow into a non-symplectic map in this context, but it is not important since the actual physical system we are integrating is not a conservative Hamiltonian system. The symplectic structure encoded in the network still seems to improve the approximation accuracy, as we show in the numerical experiments Section 4.
@@
Notice that in the figures in this paper the letters $p_{a}$ and $p_{b}$ denote the momenta. We emphasize that these momenta are the non-conservative ones, since they are the result of solving the augmented system. However, because of the projection, they coincide with the conservative ones, i.e., $\pi_{a}=p_{a}$ and $-\pi_{b}=p_{b}$, hence why we use this convention as discussed above.
Appendix B Derivation of the long-time Hamiltonian system
Let us consider the function
and evaluate its derivative in time. We evaluate the following limit
$\dot{\psi}(t,x_{0})=\lim_{h\to 0}\frac{\psi(t+h,x_{0})-\psi(t,x_{0})}{h}$
$=\mathbf{J}\nabla\mathcal{H}(\bar{\psi})(t-c\Delta t,\psi(t,x_{0}))$
as written in Equation 13.
Appendix C Proof of the a-posteriori result
Proof of 2.
To start, we notice that
$\mathcal{H}(\psi)(t,x)$ $=\mathcal{H}(\bar{\psi})(t-\Delta t\lfloor t/\Delta t\rfloor,x)$
$=H(x)+\delta(t-\Delta t\lfloor t/\Delta t\rfloor,x)$
and since $\mathbf{J}$ preserves the $\ell^{2}$ norm, we get that $\|\nabla\delta(t,x)\|_{2}\leq\varepsilon_{2}$ too. This allows us to conclude that
where $C$ is an upper bound of the Lipschitz constant of $H$ over $\Omega$. Thus,
$\left|H(\psi_{t}(x))-H(x)\right|\leq c_{2}\varepsilon_{2}t.$
Moving to the variation of $\mathcal{H}(\psi)$, we notice that
$\mathcal{H}(\psi)(t,\psi_{t}(x))-\mathcal{H}(\psi)(0,x)$ $=\mathcal{H}(\psi)(t,\psi_{t}(x))-H(\psi_{t}(x))$
$+H(\psi_{t}(x))-H(x)$
$+H(x)-\mathcal{H}(\psi)(0,x),$
which allows us to conclude
Appendix D Proof of 1
This appendix provides a proof of 1, i.e., that any Hamiltonian flow can be approximated to arbitrary accuracy, uniformly on a compact time interval $[0,\Delta t]$, by a suitable SympFlow. Our proof builds on the derivations in [56].
@@ Appendix D Proof of 1
Let $H:\mathbb{R}\times\Omega\to\mathbb{R}$ be a $C^{1}$ function, for $\Omega\subset\mathbb{R}^{2d}$ compact representing the phase space, while the first parameter denotes time. For any $\varepsilon>0$, there is a function $\widetilde{H}:\mathbb{R}\times\Omega\to\mathbb{R}$ polynomial in the phase-space variable $x\in\Omega$ and with coefficients continuously depending on time, whose exact flow $\phi_{\widetilde{H},t}$ approximates $\phi_{H,t}$ to accuracy $\varepsilon$ on $\Omega\times[0,\Delta t]$.
@@
The function $H$ is a continuous function of the $12d+$ 1 variables $(x,t)\in\Omega\times[0,\Delta t]$. Hence, it can be approximated to arbitrary accuracy on the compact set $\Omega\times[0,\Delta t]$ by a polynomial in $(x,t)$. The approximating polynomial can be chosen to also approximate the gradient of $H$ with the same accuracy. For a proof of this result see [38, Proposition 4.2].
Let us consider now a polynomial function $\widetilde{H}$ giving
@@
$\sup_{\begin{subarray}{c}t\in[0,\Delta t]\\ x\in\Omega\end{subarray}}\left\|\nabla H(t,x)-\nabla\widetilde{H}(t,x)\right\|_{\infty}<\widetilde{\varepsilon}.$
Then, we have that
$\left\|\phi_{H,t}(x)-\phi_{\widetilde{H},t}(x)\right\|_{\infty}=\left\|\mathbf{J}\int_{0}^{t}\left(\nabla H(s,\phi_{H,s}(x))-\nabla\widetilde{H}(s,\phi_{\widetilde{H},s}(x))\right)\mathrm{d}s\right\|_{\infty}=$
$=\left\|\mathbf{J}\int_{0}^{t}\left(\nabla H(s,\phi_{H,s}(x))-\nabla H(s,\phi_{\widetilde{H},s}(x))+\nabla H(s,\phi_{\widetilde{H},s}(x))-\nabla\widetilde{H}(s,\phi_{\widetilde{H},s}(x))\right)\mathrm{d}s\right\|_{\infty}$
$\leq\left(\sup_{\begin{subarray}{c}s\in[0,\Delta t]\\ x\in\Omega\end{subarray}}\left\|\nabla^{2}H(s,x)\right\|_{\infty}\right)\int_{0}^{t}\left\|\phi_{H,t}(x)-\phi_{\widetilde{H},t}(x)\right\|_{\infty}\mathrm{d}s+\widetilde{\varepsilon}\Delta t.$
Grownall’s inequality allows us to conclude that
@@
$\left\|\phi_{H,t}(x)-\phi_{\widetilde{H},t}(x)\right\|_{\infty}$ $\leq\widetilde{\varepsilon}\Delta t\exp\left(t\sup_{\begin{subarray}{c}s\in[0,\Delta t]\\ x\in\Omega\end{subarray}}\left\|\nabla^{2}H(s,x)\right\|_{\infty}\right)$
@@
$\leq\widetilde{\varepsilon}\Delta t\exp\left(\Delta t\sup_{\begin{subarray}{c}s\in[0,\Delta t]\\ x\in\Omega\end{subarray}}\left\|\nabla^{2}H(s,x)\right\|_{\infty}\right)=:\widetilde{\varepsilon}\Delta t\exp\left(\Delta tL\right).$
The lemma is proven by taking $\widetilde{\varepsilon}=\varepsilon/(\Delta t\exp(\Delta t\,L))$, where $L$ is a bound for the Lipschitz constant of $\nabla H(s,\cdot)$ uniform in $s\in[0,\Delta t]$. ∎
@@
Thanks to Lemma 1, we can now assume to work with $\widetilde{H}:\mathbb{R}\times\Omega\to\mathbb{R}$ which is a polynomial in the phase-space variable $x$ and with coefficients continuously depending on time. We can then split the flow $\phi_{\widetilde{H},t}$ into $N$ substeps of size $t/N$ as follows:
@@
In Equation 40, we denote with $\widetilde{\phi}_{nt/N}(x_{0})$ the exact solution at time $(n+1)t/N$ of the initial value problem
@@
The goal is now getting closer to the form of differential equations defining the layers of the SympFlow. To do so, we work with [56, Lemma 1], which we now state adapting the notation to ours.
@@
Let $\widetilde{H}:\mathbb{R}\times\Omega\to\mathbb{R}$ be polynomial in the phase-space variable $x\in\Omega$ and depending continuously on the first variable. There exists a function $V:\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$ polynomial in the second variable and with coefficients depending continuously on the first variable and a set of $d$ integers $\omega_{1},\ldots,\omega_{d}$, such that the Hamiltonian
$\widehat{H}(s,x)=\frac{1}{2}\left(\left\|p\right\|_{2}^{2}+2\pi q^{\top}\operatorname{diag}(\omega_{1}^{2},\ldots,\omega_{d}^{2})q\right)+\frac{t}{N}V(s,q),\,\,x=(q,p)\in\mathbb{R}^{2d},$ (41)
whose exact time-$1$ flow $\phi_{\widehat{H},1}$ is $\mathcal{O}(t^{2}/N^{2})$ close to $\widetilde{\phi}_{nt/N}$.
We do not specify it in the notation since it is already quite heavy, but this approximation depends on the time instant $nt/N$, so $\widehat{H}$ would be $\widehat{H}_{n}$. We remark that the differential equation defined by $\widehat{H}$ in Equation 41 is a second-order differential equation of the form
$\frac{\mathrm{d}^{2}}{\mathrm{d}s^{2}}q(s)=-2\pi\operatorname{diag}(\omega_{1}^{2},\ldots,\omega_{d}^{2})q(s)-\frac{t}{N}\nabla_{q}V(s,q(s)).$ (42)
For this reason, the proof in [56] is based on a perturbative argument studying Equation 42 as a perturbation of the linear oscillator obtained by removing the non-linear potential $V$.
An equivalent way to write Equation 42 is with the first-order differential equation
$\begin{bmatrix}\dot{q}(s)\\ \dot{p}(s)\end{bmatrix}=\begin{bmatrix}p(s)\\ -2\pi\operatorname{diag}(\omega_{1}^{2},\ldots,\omega_{d}^{2})q(s)-\frac{t}{N}\nabla_{q}V(s,q(s))\end{bmatrix}.$ (43)
@@
Let $\widetilde{H}:\mathbb{R}\times\Omega\to\mathbb{R}$ be polynomial in the phase-space variable $x\in\Omega$ and depending continuously on the first variable. For every $\varepsilon>0$, there exists a pair of functions $\bar{V}_{\mathrm{q}},\bar{V}_{\mathrm{p}}\in\mathcal{F}$, with
={×$\mathcal{F}=\left\{\mathbb{R}\times\mathbb{R}^{d}\ni(s,q)\mapsto w^{\top}\sigma\left(Aq+bs+c\right)\in\mathbb{R}:\,\,A\in\mathbb{R}^{h\times d},b,c,w,\in\mathbb{R}^{h},\,\,h\in\mathbb{N}\right\},$
$\sigma:\mathbb{R}\to\mathbb{R}$ continuously differentiable and not a polynomial, such that defined the Hamiltonian
$\bar{H}(s,x)=\bar{V}_{\mathrm{p}}(s,p)+\bar{V}_{\mathrm{q}}(s,q),\,\,x=(q,p)\in\mathbb{R}^{2d},$ (44)
the exact time-$1$ flow $\phi_{\bar{H},1}$ is $\varepsilon$ close to $\widetilde{\phi}_{nt/N}$.
@@
Lemma 2 is fundamental to prove Corollary 1 since we need to get a separable Hamiltonian, so that we can relate these maps with a SympFlow. Jumping over Lemma 2 would have required a significantly new analysis to directly obtain Corollary 1.
Proof.
By [44, Theorem 4.1], for any $\widetilde{\varepsilon}>0$ there are two functions $\bar{V}_{\mathrm{q}}=\bar{V}_{\mathrm{q}}(s,q)$ and $\bar{V}_{\mathrm{p}}=\bar{V}_{\mathrm{p}}(s,p)$ that belong to $\mathcal{F}$ and such that
@@
$\sup_{\begin{subarray}{c}s\in[0,1]\\ x\in\Omega\end{subarray}}\left\|\nabla\widehat{H}(s,x)-\nabla\bar{H}(s,x)\right\|_{\infty}<\widetilde{\varepsilon},$

### v2

@@ Appendix D Background on symplectic maps and Hamiltonian dynamics
Appendix E The classical mechanics of non-conservative systems: Damped Harmonic Oscillator
@@
where $L$ describes the conservative contribution to the full non-conservative Lagrangian $\Lambda$, and $K$ is a function that couples the variables together and accounts for all the non-conservative contribution to the system dynamics. Notice that, in conservative Hamiltonian systems, the function $K$ would vanish. Although, $K$ also vanishes in the physical limit (PL) [25], that is, in the linear subspace $\{(q_{a},q_{b},p_{a},p_{b})\in\mathbb{R}^{4d}:\,\,q_{a}=q_{b},\text{ and }p_{a}=p_{b}\}$, its derivatives do not necessarily vanish in the physical limit. This is important, since the derivatives of $K$ model the dissipative behavior of the system, as we will see below.
The stationarity of the action in Equation˜32 under the variations $q_{I}(t,\epsilon)=q_{I}(t,0)+\epsilon\eta_{I}(t)$ , that is $\mathrm{d}S[q_{I}]/\mathrm{d}\epsilon]_{\epsilon=0}=0$ for all $\eta_{I}(t_{i})=\eta_{I}(t_{f})=0$, $I=\{a,b\}$, corresponds to the conditions
@@
$={}$ $\int_{t_{i}}^{t_{f}}\bigg\{\eta_{a}\cdot\bigg[\frac{\partial\Lambda}{\partial q_{a}}-\frac{\mathrm{d}\pi_{a}}{\mathrm{d}t}\bigg]_{\epsilon=0}+\eta_{b}\cdot\bigg[\frac{\partial\Lambda}{\partial q_{b}}+\frac{\mathrm{d}\pi_{b}}{\mathrm{d}t}\bigg]_{\epsilon=0}\bigg\}\,\mathrm{d}t$
$+{}$ $\bigg[\eta_{a}(t)\cdot\pi_{b}(t)-\eta_{b}(t)\cdot\pi_{a}(t)\bigg]_{t=t_{i}}^{t_{f}}$ (33)
@@
We remark that $\pi_{a}=-\pi_{b}$ can be inferred by noticing that $\pi_{a}+\pi_{b}=p_{a}-p_{b}-\lambda(q_{a}-q_{b})=0$, where $p_{a}$ and $p_{b}$ are the conservative momenta derived from the conservative Lagrangian as $p_{I}=\partial_{\dot{q}_{I}}L(q_{I},\dot{q}_{I})$, $I\in\{a,b\}$.
@@
$\begin{bmatrix}\dot{q}_{a}(t)\\ \dot{q}_{b}(t)\\ \dot{p}_{a}(t)\\ \dot{p}_{b}(t)\end{bmatrix}=\begin{bmatrix}\frac{\pi_{a}(t)}{m}+\frac{\lambda}{2m}(q_{a}(t)-q_{b}(t))\\ -\frac{\pi_{b}(t)}{m}-\frac{\lambda}{2m}(q_{a}(t)-q_{b}(t))\\ -\frac{\lambda}{2m}(\pi_{a}(t)-\pi_{b}(t))-kq_{a}(t)\\ \frac{\lambda}{2m}(\pi_{a}(t)-\pi_{b}(t))+kq_{b}(t)\end{bmatrix}$. (40)
Notice that the physical limit in Equation˜36 is an invariant submanifold of $\mathbb{R}^{4d}$ with respect to the dynamics, since $\dot{\pi}_{a}+\dot{\pi}_{b}=0$ whenever $q_{a}=q_{b}$, and also $\dot{q}_{a}-\dot{q}_{b}=0$. To ensure that our neural networks preserve this subspace as well, we apply a projection step at the end of the network, which projects over the subspace in Equation˜36 and is defined as
@@
We remark that this projection turns the SympFlow into a non-symplectic map in this context, but it is not important since the actual physical system we are integrating is not a conservative Hamiltonian system. The symplectic structure encoded in the network still seems to improve the approximation accuracy, as we show in the numerical experiments Section˜4.
@@
Notice that in the figures in this paper the letters $p_{a}$ and $p_{b}$ denote the momenta. We emphasize that these momenta are the non-conservative ones, since they are the result of solving the augmented system. However, because of the projection, they coincide with the conservative ones, i.e., $\pi_{a}=p_{a}$ and $\pi_{b}=p_{b}$, hence why we use this convention as discussed above.
Appendix F Proof of Theorem 1
This appendix provides a proof of ˜1, i.e., that any Hamiltonian flow can be approximated to arbitrary accuracy, uniformly on a compact time interval $0,\Delta t]$, by a suitable SympFlow. Our proof builds on the derivations in [64].
@@ Appendix F Proof of Theorem 1
Let $H:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}$ be a twice-continuously differentiable function. Fix $\Delta t>0$. Consider a compact set $\Omega\subset\mathbb{R}^{2d}$ and assume $\Omega$ is forward invariant for $\phi_{H,t}$. For any $\varepsilon>0$, there is a function $\widetilde{H}:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}$ polynomial in the phase-space variable $x\in\mathbb{R}^{2d}$ and with coefficients continuously depending on time, whose exact flow $\phi_{\widetilde{H},t}$ approximates $\phi_{H,t}$ to accuracy $\varepsilon$ on $\Omega\times[0,\Delta t]$.
@@
Define
$\Omega_{\varepsilon}:=\Omega+B_{\infty}(\varepsilon)=\left\{x\in\mathbb{R}^{2d}:\min_{y\in\Omega}\|x-y\|_{\infty}\leq\varepsilon\right\}.$
Call $\Omega_{\varepsilon,c}$ a compact and convex set containing $\Omega_{\varepsilon}$, and define
@@
$L:=\sup_{\begin{subarray}{c}s\in[0,\Delta t]\\ x\in\Omega_{\varepsilon,c}\end{subarray}}\left\|\nabla^{2}_{x}H(s,x)\right\|_{\infty},\,\text{ and }\,\widetilde{\varepsilon}:=\frac{\varepsilon}{\Delta t\exp(\Delta tL)},$
where $\nabla_{x}^{2}H$ is the Hessian in the second variable. $L$ is finite as $\Omega_{\varepsilon,c}$ is compact and $H$ is twice-continuously differentiable on the whole space. Consider a polynomial function $\widetilde{H}:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}$ such that
@@
$\sup_{\begin{subarray}{c}t\in[0,\Delta t]\\ x\in\Omega_{\varepsilon}\end{subarray}}\left|H(t,x)-\widetilde{H}(t,x)\right|<\widetilde{\varepsilon}$, (45)
@@
$\sup_{\begin{subarray}{c}t\in[0,\Delta t]\\ x\in\Omega_{\varepsilon}\end{subarray}}\left\|\nabla_{x}H(t,x)-\nabla_{x}\widetilde{H}(t,x)\right\|_{\infty}<\widetilde{\varepsilon}$, (46)
whose existence is guaranteed by the Weierstrass Approximation Theorem [47, Theorem 1.6.2] since $\Omega_{\varepsilon}$ is compact. We remark that $\nabla_{x}H$ stands for the gradient of $H$ in the second variable.
Since the approximation only holds on $\Omega_{\varepsilon}$, we introduce
$\tau(x):=\min\left\{\Delta t,\inf\{t\in[0,+\infty):\,\phi_{\widetilde{H},t}(x)\in\partial\Omega_{\varepsilon}\}\right\}.$
Since $\widetilde{H}$ is polynomial, the associated vector field is $C^{1}$ (hence locally Lipschitz in the state variable). Therefore, by the Picard–Lindelöf Theorem, for every $x\in\Omega$ there exists a unique local solution $t\mapsto\phi_{\widetilde{H},t}(x)$, which is in particular continuous and satisfies $\phi_{\widetilde{H},0}(x)=x\in\Omega\subset\Omega_{\varepsilon}$. Consequently $\tau(x)>0$. Moreover, define
$B:=\sup_{\begin{subarray}{c}t\in[0,\Delta t]\\ z\in\Omega_{\varepsilon,c}\end{subarray}}\left\|\nabla_{x}\widetilde{H}(t,z)\right\|_{\infty}<\infty,$
which is finite by continuity of $\nabla_{x}\widetilde{H}$ on the compact set $0,\Delta t]\times\Omega_{\varepsilon,c}$. Then, for every $x\in\Omega$ and $t\in[0,\tau(x)]$, we have
$\left\|x-\phi_{\widetilde{H},t}(x)\right\|_{\infty}\leq\int_{0}^{t}\left\|\nabla_{x}\widetilde{H}(s,\phi_{\widetilde{H},s}(x))\right\|_{\infty}\,\mathrm{d}s\leq Bt,$
since $\phi_{\widetilde{H},s}(x)\in\Omega_{\varepsilon}\subset\Omega_{\varepsilon,c}$ for all $s\in[0,t]$. If $\tau(x)<\min\{\Delta t,\varepsilon/B\}$, then
$\min_{y\in\Omega}\|y-\phi_{\widetilde{H},\tau(x)}(x)\|_{\infty}\leq\left\|x-\phi_{\widetilde{H},\tau(x)}(x)\right\|_{\infty}\leq B\tau(x)<\varepsilon,$
contradicting the definition of $\tau(x)$. Hence $\tau(x)\geq\min\{\Delta t,\varepsilon/B\}>0$ for every $x\in\Omega$.
For any $x\in\Omega$ and $t\in[0,\tau(x)]$, we can write
$\left\|\phi_{H,t}(x)-\phi_{\widetilde{H},t}(x)\right\|_{\infty}=\left\|\mathbf{J}\int_{0}^{t}\left(\nabla_{x}H(s,\phi_{H,s}(x))-\nabla_{x}\widetilde{H}(s,\phi_{\widetilde{H},s}(x))\right)\mathrm{d}s\right\|_{\infty}=$
$=\left\|\mathbf{J}\int_{0}^{t}\left(\nabla_{x}H(s,\phi_{H,s}(x))-\nabla_{x}H(s,\phi_{\widetilde{H},s}(x))+\nabla_{x}H(s,\phi_{\widetilde{H},s}(x))-\nabla_{x}\widetilde{H}(s,\phi_{\widetilde{H},s}(x))\right)\mathrm{d}s\right\|_{\infty}$
$\leq\underbrace{L\int_{0}^{t}\left\|\phi_{H,s}(x)-\phi_{\widetilde{H},s}(x)\right\|_{\infty}\mathrm{d}s}_{(a)}+\int_{0}^{t}\left\|\nabla_{x}H(s,\phi_{\widetilde{H},s}(x))-\nabla_{x}\widetilde{H}(s,\phi_{\widetilde{H},s}(x))\right\|_{\infty}\mathrm{d}s.$
$(a)$ follows from the Lipschitz regularity of the gradient of twice-continuously differentiable functions on compact and convex sets. Since $t\in[0,\tau(x)]$, then Equation˜46 applies and we get
$\left\|\phi_{H,t}(x)-\phi_{\widetilde{H},t}(x)\right\|_{\infty}\leq L\int_{0}^{t}\left\|\phi_{H,s}(x)-\phi_{\widetilde{H},s}(x)\right\|_{\infty}\mathrm{d}s+\widetilde{\varepsilon}t.$
Gronwall’s inequality allows us to conclude that, for any $x\in\Omega$ and $t\in[0,\tau(x)]$ we can write
$\left\|\phi_{H,t}(x)-\phi_{\widetilde{H},t}(x)\right\|_{\infty}\leq\widetilde{\varepsilon}t\exp\left(Lt\right)\leq\widetilde{\varepsilon}\Delta t\exp\left(L\Delta t\right)=\varepsilon.$
We will now show that $\tau(x)=\Delta t$. Assume, by contradiction, that $\tau(x)<\Delta t$ for an $x\in\Omega$, i.e., $\phi_{\widetilde{H},\tau(x)}(x)\in\partial\Omega_{\varepsilon}$. Then, by the forward invariance of $\Omega$, $\phi_{{H},\tau(x)}(x)\in\Omega$ and hence
$\|\phi_{{H},\tau(x)}(x)-\phi_{\widetilde{H},\tau(x)}(x)\|_{\infty}\geq\min_{y\in\Omega}\|y-\phi_{\widetilde{H},\tau(x)}(x)\|_{\infty}=\varepsilon.$
This is a contradiction since the previous derivation implies
$\|\phi_{{H},\tau(x)}(x)-\phi_{\widetilde{H},\tau(x)}(x)\|_{\infty}\leq\widetilde{\varepsilon}\tau(x)\exp\left(L\tau(x)\right)<\widetilde{\varepsilon}\Delta t\exp\left(L\Delta t\right)=\varepsilon.$
We conclude that $\tau(x)=\Delta t$ for every $x$ and the desired result follows. ∎
@@
Thanks to Lemma˜2, we can now assume to work with $\widetilde{H}:\mathbb{R}\times\Omega\to\mathbb{R}$ which is a polynomial in the phase-space variable $x$ and with coefficients continuously depending on time. We can then split the flow $\phi_{\widetilde{H},t}$ into $N$ substeps of size $t/N$ as follows:
@@
In Equation˜47, we denote with $\widetilde{\phi}_{nt/N}(x_{0})$ the exact solution at time $(n+1)t/N$ of the initial value problem
@@
The goal is now getting closer to the form of differential equations defining the layers of the SympFlow. To do so, we work with [64, Lemma 1], which we now state adapting the notation to ours.
@@
Let $\widetilde{H}:\mathbb{R}\times{\mathbb{R}^{2d}}\to\mathbb{R}$ be polynomial in the phase-space variable $x\in{\mathbb{R}^{2d}}$ and depending continuously on the first variable. Fix $t\in[0,+\infty)$ and let $\Omega\subset\mathbb{R}^{2d}$ be an arbitrary compact set. There exists a function $V:\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$ polynomial in the second variable and with coefficients depending continuously on the first variable and a set of $d$ integers $\omega_{1},\ldots,\omega_{d}$, such that the Hamiltonian
$\widehat{H}(s,x)=\pi\left(\left\|p\right\|_{2}^{2}+q^{\top}\operatorname{diag}(\omega_{1}^{2},\ldots,\omega_{d}^{2})q\right)+\frac{t}{N}V(s,q),\,\,x=(q,p)\in\mathbb{R}^{2d},$ (48)
with time-$1$ flow $\phi_{\widehat{H},1}$ which is $\mathcal{O}(t^{2}/N^{2})$ close to $\widetilde{\phi}_{nt/N}$ in $\Omega$, i.e.,
$\sup_{x\in\Omega}\left\|\phi_{\widehat{H},1}(x)-\widetilde{\phi}_{nt/N}(x)\right\|_{\infty}=\mathcal{O}(t^{2}/N^{2}).$
We remark that the big-O notation $f(t)=\mathcal{O}(g(t))$, for a pair of functions $f,g:[0,+\infty)\to\mathbb{R}$ means that there is a constant $c>0$ such that $|f(t)|\leq c|g(t)|$ for all $t\geq 0$. We do not specify it in the notation since it is already quite heavy, but this approximation depends on the time instant $nt/N$, so $\widehat{H}$ would be $\widehat{H}_{n}$. We remark that the differential equation defined by $\widehat{H}$ in Equation˜48,
$\begin{bmatrix}\dot{q}(s)\\ \dot{p}(s)\end{bmatrix}=\begin{bmatrix}2\pi p(s)\\ -2\pi\operatorname{diag}(\omega_{1}^{2},\ldots,\omega_{d}^{2})q(s)-\frac{t}{N}\nabla_{q}V(s,q(s))\end{bmatrix}$, (49)
can be seen as a perturbation of the linear oscillator
$\frac{\mathrm{d}^{2}}{\mathrm{d}s^{2}}q(s)=-4\pi^{2}\operatorname{diag}(\omega_{1}^{2},\ldots,\omega_{d}^{2})q(s).$ (50)
@@
Let $\widetilde{H}:\mathbb{R}\times{\mathbb{R}^{2d}}\to\mathbb{R}$ be polynomial in the phase-space variable $x\in{\mathbb{R}^{2d}}$ and depending continuously on the first variable. Consider an arbitrary $\varepsilon>0$, $t\in[0,+\infty)$, and a compact set $\Omega\subset\mathbb{R}^{2d}$. Let
$\mathcal{F}{{}_{\sigma}}=\left\{\mathbb{R}\times\mathbb{R}^{d}\ni(s,q)\mapsto w^{\top}\sigma\left(Aq+bs+c\right)\in\mathbb{R}:\,\,A\in\mathbb{R}^{h\times d},b,c,w,\in\mathbb{R}^{h},\,\,h\in\mathbb{N}\right\},$ (51)
with $\sigma:\mathbb{R}\to\mathbb{R}$ continuously differentiable and not a polynomial, which is applied entrywise. There exists a Hamiltonian function
$\bar{H}(s,x)=\bar{V}_{\mathrm{p}}(s,p)+\bar{V}_{\mathrm{q}}(s,q),\,\,\bar{V}_{\mathrm{p}},\,\bar{V}_{\mathrm{q}}\in\mathcal{F}_{\sigma},x=(q,p)\in\mathbb{R}^{2d},$ (52)
whose exact time-$1$ flow $\phi_{\bar{H},1}$ is $\mathcal{O}(t^{2}/N^{2})$ close to $\widetilde{\phi}_{nt/N}$, i.e.,
$\sup_{x\in\Omega}\left\|\widetilde{\phi}_{nt/N}(x)-\phi_{\bar{H},1}(x)\right\|_{\infty}=\mathcal{O}(t^{2}/N^{2}).$
@@
Lemma˜3 is fundamental to prove Corollary˜1 since we first need to get a separable Hamiltonian, so that we can relate these maps with a SympFlow. Jumping over Lemma˜3 would have required a significantly new analysis to directly obtain Corollary˜1.
Proof of Corollary˜1.
This proof mimics the one of Lemma 2. Lemma 3 ensures that there exists $\widehat{H}$ as in (48), and a constant $c>0$, such that
$\sup_{x\in\Omega}\left\|\phi_{\widehat{H},1}(x)-\widetilde{\phi}_{nt/N}(x)\right\|_{\infty}\leq c\frac{t^{2}}{N^{2}}.$
Let us define
$\widehat{\Omega}:=\left\{y\in\mathbb{R}^{2d}:\,\,y=\phi_{\widehat{H},s}(x),\,x\in\Omega,\,s\in[0,1]\right\},$
which is compact. Let $\widehat{\Omega}_{\varepsilon}=\widehat{\Omega}+B_{\infty}(\varepsilon)$, and call $\widehat{\Omega}_{\varepsilon,c}\subset\mathbb{R}^{2d}$ a compact and convex set containing $\widehat{\Omega}_{\varepsilon}$. Set
@@
$L:=\sup_{\begin{subarray}{c}s\in[0,1]\\ x\in\widehat{\Omega}_{\varepsilon,c}\end{subarray}}\left\|\nabla_{x}^{2}\widehat{H}(s,x)\right\|_{\infty},\,\,\widehat{\varepsilon}:=\frac{\varepsilon}{\exp(L)}.$
By [51, Theorem 4.1], there exists a pair of functions $\bar{V}_{\mathrm{q}}=\bar{V}_{\mathrm{q}}(s,q)$ and $\bar{V}_{\mathrm{p}}=\bar{V}_{\mathrm{p}}(s,p)$ that belong to $\mathcal{F}{{}_{\sigma}}$ and such that
$\sup_{\begin{subarray}{c}s\in[0,1]\\ x\in\widehat{\Omega}_{\varepsilon}\end{subarray}}\left|\widehat{H}(s,x)-\bar{H}(s,x)\right|<\widehat{\varepsilon},\,\,\sup_{\begin{subarray}{c}s\in[0,1]\\ x\in\widehat{\Omega}_{\varepsilon}\end{subarray}}\left\|\nabla\widehat{H}(s,x)-\nabla\bar{H}(s,x)\right\|_{\infty}<\widehat{\varepsilon}$, (53)